TIME-RESOLVED FOURIER TRANSFORM INFRARED EMISSION SPECTRA OF HNC/HCN

\underline{K. KAWAGUCHI}, A. FUJIMOTO, Department of Chemistry, Faculty of Science, Okayama University, Okayama 700-8530, Japan.

A continuously scanning FT spectrometer (Bruker IFS 120 HR) has been applied for observing time-resolved spectra after pulse phenomena by a multi-sampling method with a help of a chip computer SX^{a}. In the present study, we used a pulse discharge with a duration of 40 μsec in a mixture of N_{2} (120 mTorr) and CH_{4} (15 mTorr) to produce HNC and HCN emissions which were monitored by the ν_{1} bands (C-H stretch in HCN or NH stretch in HNC) with a resolution of 0.06 cm^{−1}. The cell was cooled down by liquid nitrogen or dryice. Thirty data were collected to produce 30 interferograms with a 2 μsec interval. The emission intensities from v_{1} = 1 of HCN and HNC decreased for 15 μsec after turning off the discharge and increased later. The emission from v_{1} = 2 of HCN showed similar time profile as the v_{1} = 1 emission. However, the emission from v_{2}=2 of HNC showed a different time profile.

aK. Kawaguchi, O. Baskakov, Y. Hosaki, Y. Hama, and C. Kugimiya, Chem. Phys. Lett. 369, 293(2003)